the freight train took 10 hours to make the same trip the express train made in 6 hours. Find the speed of the express train if it was 20 mph faster than the freight train
Answer by josgarithmetic(39616)

The trip distance is d.
Freight train's speed is r.

______________speed________time________distance
freight_______r_____________10_________d=r*10
express_______r+20___________6_________d=(r+20)*6

EQUAL DISTANCE TRIP!